When density and molarity is given and ask to find the molality what is the formula to be used

Grade:12th pass

1 Answers

Deepak Kumar
6 years ago
Formula ism = (1000*M)/[(1000*d) - (M * MM) ]where,m = required molalityM = given molarityd = density of solution in g/cm^3MM = molar mass in gramsDerivation of formulaLet M is given molarity i.e. there is M number of moles of solute in 1 litre of solutiond = density of solution in g/ccMass of solution = d * 1000 SoMass of solvent = (1000*d) - (M*MM)Now,Molality = no. Of moles of solute/ mass of solvent in kgSo, Molality m = (1000*M)/[(1000*d) - (M * MM) ]